OpenStudy (anonymous):

How would you graph -5/4 < x?

OpenStudy (anonymous):

\[ -\dfrac {5}{4} < x \rightarrow x > - \dfrac {5}{4}. \]Draw the graph of x = 5/4 and shade the region above.

OpenStudy (anonymous):

Thats what I need help with. The actual graph.

OpenStudy (anonymous):

With the graph of what? \[ x = - \dfrac {5}{4}? \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Hint: It's vertical.

OpenStudy (anonymous):

Well, i know that. I just need to know what numbers go in between them.

OpenStudy (anonymous):

x has a constant value. y varies over all reals.